OpenStudy (anonymous):

Is anyone hear good with linear functions? If so then please help me with this one... 1) A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 40 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. A f(x) = - 2x + 40 B f(x) = 2x + 40 C f(x) = - 2x + 50 D f(x) = 2x + 50

OpenStudy (anonymous):

lol I meant here not hear

OpenStudy (anonymous):

I think its C but Im not quite sure...

OpenStudy (anonymous):

yeah you are correct.. its C..

OpenStudy (anonymous):

Yes. it is C

OpenStudy (anonymous):

Sweet thanks... also can anyone help with another question?

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. A: y - 6 = one half(x + 3) B: y = one halfx - 6 C: y + 3 = one half(x - 6) D: x - 2y = 12

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yeah its B..

OpenStudy (anonymous):

Oh... I messed up somewhere and got D but thanks again for the help! :D

OpenStudy (anonymous):

No Problem:)